Two asteroids of equal mass in the asteroid belt between mars and Jupiter collide with a glancing blow. Asteroid A, which was initially travelling at 40m/s is deflected 30 degrees from its original position while Asteroid B travels at 45 degrees to the original direction of A. Find the speed of each Asteroid after collision and what fraction of the original kinetic energy of Asteroid A dissipates during this collision
1
Expert's answer
2020-07-13T11:45:41-0400


According to the law of conservation of momentum, for the initial and final momenta we can write

"\vec{p}_A+\vec{p}_b=\vec{p'}_A+\vec{p'}_B."

If the asteroid B was initially at rest, its momentum is zero, and the equation becomes


"\vec{p}_A=\vec{p'}_A+\vec{p'}_B."


Conservation of momentum along x- and y-axes:


"Ox: mv_A=mv'_A\text{ cos}30°+mv_B'\text{ cos}45°,\\
Oy: 0=mv_A'\text{ sin}30°-mv_B'\text{ sin}45°."

The equation for y-axis gives us a relation between final speeds of A and B:


"v'_A\text{ sin}30°=v'_B\text{ sin}45°,\\
v_A'=v'_B\sqrt2."

Therefore, if we substitute this into the equation for x-axis, we will get the final speed of A:


"v_A=v'_B\sqrt2\text{ cos}30°+v'_B\text{ cos}45°,\\
v'_B=\frac{v_A}{\sqrt2\text{ cos}30°+\text{ cos}45°},\\\space\\
v'_A=\frac{40}{\text{ cos}30°+\sqrt2\text{ cos}45°}=20.7\text{ m/s}.\\\space\\
v'_A=v'_B\sqrt2=29.3\text{ m/s}."

To find what fraction of the original kinetic energy of asteroid A dissipates during this collision, determine the initial kinetic energy of A:


"K_i=\frac{mv_A^2}{2}."

The final energy of A was


"K_f=\frac{m{v'_A}^2}{2}."

Energy dissipated during the collision is


"\Delta K=K_i-K_f."

It corresponds to the following fraction of the original kinetic energy:


"\epsilon=\frac{\Delta K}{K_i}=1-\frac{v_A'^2}{v_A^2}=0.46."

About 46% of original energy dissipated.